# Stability Regularized Cross-Validation

### v1

@@ 2.2 Generalization Bound
By combining our definitions and notation, letting $M$ represent an upper bound on the loss $\ell(\bm{\beta}(\bm{\theta},\bm{x}_{i}),y_{i})$ for any model $\bm{\beta}(\bm{\theta})$ and any datapoint $(\bm{x}_{i},y_{i})$ (e.g., if $(\bm{x}_{i},y_{i})$ are drawn from a bounded domain), and letting $\mathcal{S}$ denote a test set of observations drawn from the same distribution as our training set (but not seen by the model), the following result follows from Chebyshev’s inequality (proof deferred to Section B.1):
Theorem 1.
@@
Suppose the training data $(\bm{x}_{i},y_{i})_{i\in[n]}$ are drawn from an unknown distribution $\mathcal{D}$ such that $M$ and $\mu_{h}$ are finite constants. Further, suppose $n$ is exactly divisible by $k$ and each $\mathcal{N}_{j}$ is of cardinality $n/k$. Then, the following bound on the test error holds with probability at least $1-\Omega$:
$\frac{1}{|\mathcal{S}|}\sum_{i\in\mathcal{S}}\ell\left(y_{i},\beta(\bm{\theta},\bm{x}_{i})\right)\leq\frac{1}{n}\sum_{j\in[k]}h_{j}(\bm{\theta})+\sqrt{\frac{M^{2}+6Mk\mu_{h}}{2k\Omega}}.$ (6)
Remark 1 (To Train or to Validate in (6)).
@@
A similar bound to (6) can be derived using the empirical risk instead of kCV [11, Theorem 11]. However, this bound has a larger constant ($12$ instead of $6$) and still involves the expensive pointwise hypothesis stability defined by [11].
@@
Theorem 1 reveals that, if the number of folds $k$ increases with $n$, $M$ is finite, and the hypothesis stability $\mu_{h}$ decreases with $n$, then the kCV error generalizes to the test set with high probability as $n$ becomes large. Moreover, when models have the same cross-validation error, hypothesis stability, and loss bound $M$, training on more folds results in a stronger generalization bound.
@@
We remark that Theorem 1 implicitly justifies the use of regularization in machine learning, because regularization implicitly controls the hypothesis stability $\mu_{h}$, leading to better generalization properties when $\mu_{h}$ is lower. Indeed, Bousquet and Elisseeff [11, Theorem 22] provides a result formalizing this notion in the context of supervised learning for Reproducing Kernel Hilbert Spaces.
@@
Unfortunately, in preliminary experiments, we found that Equation (6)’s bound is often excessively conservative in practice, especially when $n\gg p$. This conservatism stems from using Chebyshev’s inequality in the proof of Theorem 1, which is known to be tight for discrete measures but excessively conservative over continuous measures [5], especially unimodal continuous measures [55]. Thus, motivated by the robust optimization literature, where probabilistic guarantees are used to motivate uncertainty sets but less stringent guarantees are used in practice to avoid excessive conservatism [see 20, Section 3, for a discussion], we leverage Theorem 1 to propose a new approach to cross-validation in the rest of this section.
2.3 Stability Regularization and Nested Cross-Validation
@@
Motivated by the idea that more stable models are less likely to disappoint out-of-sample (Theorem 1), we propose selecting models that minimize a weighted sum of the kCV error and the hypothesis stability. This corresponds to selecting $\bm{\theta}$ through the optimization problem
$\bm{\theta}\in\operatorname*{arg\,min}_{\bm{\theta}\in\bm{\Theta}}\frac{1}{n}\sum_{j\in[k]}h_{j}(\bm{\theta})+\lambda\mu_{h}(\bm{\theta}).$ (7)
@@
Notably, this procedure still requires a hyperparameter $\lambda$, which the user must select. Accordingly, we invoke nested cross-validation, which has been empirically shown to be significantly less vulnerable to out-of-sample disappointment than regular cross-validation [13, 3]. We perform an outer loop over candidate $\lambda$ values. For each $\lambda$, we run an inner (nested) k-fold cross-validation: on each fold, train a model with that $\lambda$ (using the remaining folds with stability regularization) and measure its validation error. We then choose the $\lambda$ that yields the lowest average inner validation error across the $k$ folds.
@@
$\lambda\in$ $\arg\min_{\lambda\in\bm{\Lambda}}\sum_{j\in[k]}\sum_{i\in\mathcal{N}_{j}}\ell(y_{i},\bm{\beta}^{(\mathcal{N}_{j})}(\bm{\theta}_{j}^{\star},\bm{x}_{j})),$ (8)
where $\bm{\beta}^{(\mathcal{N}_{j})}(\bm{\theta}_{j}^{\star})$ denotes a model trained on all data but the fold $\mathcal{N}_{j}$ with hyperparameters $\bm{\theta}_{j}^{\star}$, and $\bm{\theta}_{j}^{\star}$ denotes an optimal solution to the following lower-level problem, which cross-validates $\bm{\theta}$ with the $j$th fold of the data omitted and $\lambda$ fixed:
$\bm{\theta}^{\star}_{j}\in\operatorname*{arg\,min}_{\bm{\theta}\in\bm{\Theta}}\frac{1}{n-|\mathcal{N}_{j}|}\sum_{l\in[k]:l\neq j}\sum_{i\in\mathcal{N}_{l}}\ell(y_{i},\bm{\beta}^{(\mathcal{N}_{j,l})}(\bm{\theta}))+\lambda\bm{\mu}_{h}(\bm{\theta}),$
where $\bm{\mu}_{h}(\bm{\theta})$ is calculated without reference to the $j$th fold of the data in this case. This corresponds to selecting $\lambda$ in a manner that ensures that models $\bm{\theta^{\star}_{j}}$ perform well on average, as in the standard nested cross-validation paradigm.
Finally, once $\lambda$ is selected, we fix $\lambda$ and select $\bm{\theta}$ by minimizing (7). This selects models that are robust to omitting one fold of the data, and tend to perform well out-of-sample as estimated by the nested cross-validation procedure, at the price of increasing the cost of hyperparameter selection. For completeness, we formalize this procedure in Algorithm 1 (see Appendix C).
We acknowledge that even nested CV errors are optimistically biased estimates (as ordinary CV errors are). However, various authors [13, 3, 4] and our experiments indicate that the nested CV error is a more accurate estimator of test performance than the standard CV error. We view our nesting approach as a practical improvement over the status quo, while leaving room for future refinements.
@@ 3 Numerical Experiments
In this section, we evaluate the numerical performance of the nested regularized cross-validation scheme proposed in Section 2. All experiments in this section were implemented in Julia version $1.9$, using Mosek version $11.0$ to solve all conic optimization problems, and conducted on a standard MacBook Pro laptop with a $36$ GB Apple M3 CPU and $36$ GB main memory.
@@ 3.1 Ridge Regularized Best Subset Selection
We now benchmark our proposed nested cross-validation scheme on sparse regression for a suite of commonly studied real-world datasets. Specifically, we benchmark a cyclic coordinate descent scheme for $\ell_{0}$$\ell_{2}^{2}$ sparse regression, where we repeatedly solve the following lower-level problem for different values of the sparsity parameter $\tau$ and the regularization hyperparameter $\gamma$
$\min_{\bm{\beta}\in\mathbb{R}^{p},\bm{z}\in\{0,1\}^{p}}\;$ $\frac{1}{2}\|\bm{X}\bm{\beta}-\bm{y}\|_{2}^{2}+\frac{1}{2\gamma}\sum_{j=1}^{p}\frac{\beta_{j}^{2}}{z_{j}}\ \text{s.t.}\ \sum_{j=1}^{p}z_{j}\leq\tau,$
using the greedy rounding algorithm described by [59, 9] to obtain near-optimal solutions in a practically tractable amount of time. In particular, we iteratively minimize the $k$-fold cross-validation error with respect to $\tau$ (with $\gamma$ fixed) and with respect to $\gamma$ (with $\tau$ fixed) until we either cycle or exceed a limit of $10$ iterations. After converging, we fit a model to the full dataset with the hyperparameters ($\tau^{\star},n_{\text{train}}\gamma^{\star}/n)$, where $n_{\text{train}}$ is the number of observations with one fold of the data left out as in [36], to account for the extra fold when fitting a model to the full dataset. Moreover, we perform the same procedure for nested $k$-fold cross-validation with stability regularization. Note that for our cyclic coordinate descent schemes, we set the largest permissible value of $\tau$ such that $\tau\log\tau\leq n$, because Gamarnik and Zadik [17, Theorem 2.5] demonstrated that, up to constant terms and under certain assumptions on the data generation process, on the order of $\tau\log\tau$ observations are necessary to recover a sparse model with binary coefficients. In preliminary experiments, we relaxed this requirement to $\tau\leq p$ and found that this did not change the optimal value of $\tau$. Moreover, we use a grid size of $20$ values of $\gamma$ log-uniformly distributed over $\{10^{-3},10^{3}\}$.
@@
  • The L0Learn.cvfit method implemented in the L0Learn R package [cf. 30], with a grid of $10$ different values of $\gamma$ and default parameters otherwise.
@@
Table 1 depicts the dimensionality of each dataset, the average $k$-fold cross-validation error (“CV”) or nested $k$-fold cross-validation error (“nCV”), the average test set error (“MSE”), and the sparsity attained by our coordinate descent scheme without any stability regularization or nesting (kCV), our cyclic coordinate descent with nested stability-adjusted cross-validation (nested-kCV), and the performance of MCP on each dataset. We also report summary statistics in terms of the average percentage improvement of the nested procedure compared to the MSE without nesting, so each dataset is weighted equally. We remark that the average (geometric mean over the average for each dataset) runtime was $14.8$ seconds for sparse regression, $592$ seconds for sparse regression with nesting, and under $1$ second for all other methods. Table 4 (Appendix E) also shows to performance of other methods from the literature (SCAD, GLMNet, and L0Learn) in the same datasets.
We measure the improvement from nested cross-validation vs. $k$-fold cross-validation by computing the average MSE across each dataset, normalizing by the MSE of $k$-fold cross-validation on that dataset to compute a percentage improvement across each dataset, and then taking the geometric mean across all datasets (to account for the fact that percentage improvement is an asymmetric measure).
@@
Avg. - - 26.75 0.056 0.096 25.80 0.089 0.099 5.90 0.105 0.085
Std. dev - - 11.62 0.072 0.136 8.92 0.121 0.134 2.45 0.098 0.098
Table 1: Average performance of methods across a suite of real-world datasets where the ground truth is unknown (and may not be sparse), sorted by how overdetermined the dataset is ($n/p$), and separated into the underdetermined and overdetermined cases. We also report standard deviations across the underdetermined and overdetermined datasets.
We observe that across the overdetermined datasets, nested cross-validation improves the out-of-sample MSE of sparse regression by $10.0\%$ on average, while across the underdetermined datasets it worsens the out-of-sample MSE by $5.92\%$ on average, for an overall improvement of $4.85\%$ on average. Moreover, the nested kCV error is, on average, $0.9\%$ smaller than the test set error for the stability-adjusted method, while the kCV error is on average $21.8\%$ smaller than the test set error for sparse ridge regression. This is especially visible for the most underdetermined datasets (Vessel and Riboflavin), where the kCV error for sparse ridge regression is significantly lower than all other methods, but this does not translate to better out-of-sample performance. In contrast, the nested kCV error is a substantially more accurate estimator of the test-set error.
@@ 3.2 Tree-Based Methods
We benchmark cyclic coordinate descent for the following two methods:
  • A Julia implementation of CART [12] via the DecisionTree.jl package, where we iteratively optimize the $5$-fold and nested $5$-fold cross-validation error with respect to the tree depth over a grid of the integers $\{1,\ldots,10\}$ and the min_samples parameter over a grid of the integers $\{2,\ldots,10\}$, with the tree depth initially fixed to $5$ for both approaches. Note that the DecisionTree.jl package holds all unspecified parameters to their default values, and thus all remaining CART parameters will take default parameters in this experiment.
  • A Julia implementation of XGBoost [14] via the XGBoost.jl package, where we iteratively optimize the $5$-fold and nested $5$-fold cross-validation error with respect to the max_depth parameter over a grid of the integers $\{1,\ldots,10\}$ and the subsample parameter over the grid [^3] We initially tried increments of $0.05$, where XGBoost selected identical hyperparameters with/without stability regularization. $\{0.01,0.02,0.03,\ldots,1.0\}$, with the tree depth initially fixed to $5$ for both approaches.
Table 2 depicts the dimensionality of each dataset, the average $k$-fold cross-validation error (“CV”) or nested $k$-fold cross-validation error (“nCV”), and the average test set error (“MSE”), with and without nested cross-validation. For both methods, the first two columns correspond to the performance of $k$-fold cross-validation, and the second two correspond to nested stability-adjusted cross-validation.
As in the previous experiment, we measure the improvement from nested cross-validation vs. $k$-fold cross-validation by computing the average MSE across each dataset, normalizing by the MSE of $k$-fold cross-validation on that dataset to compute a percentage improvement across each dataset, and then taking the geometric mean across all datasets (to account for the fact that percentage improvement is an asymmetric measure). The average (geometric mean over averages for each dataset) runtime was $5.1$s (resp. $285.8$s) seconds for XGBoost without (with) nesting, and $<0.01$s (resp. $<0.01$s) for CART with/without nesting.
@@
Wine 6497 11 0.534 0.530 0.539 0.527 0.422 0.406 0.429 0.406
Housing 506 13 19.149 20.33 20.03 20.35 12.83 12.95 14.30 12.95
Auto-MPG 392 25 14.93 17.09 16.83 17.34 10.37 11.98 11.23 11.28
Hitters 263 19 0.046 0.059 0.052 0.059 0.035 0.039 0.038 0.039
Prostate 97 8 0.691 0.803 0.780 0.814 0.545 0.672 0.609 0.667
Servo 167 19 0.314 0.403 0.390 0.394 0.18 0.163 0.217 0.163
Toxicity 38 9 0.078 0.102 0.089 0.101 0.066 0.074 0.079 0.075
Steam 25 8 1.230 1.067 1.542 1.069 0.844 1.251 0.991 1.282
Alcohol2 44 21 0.756 0.966 0.798 0.874 0.886 1.054 0.980 1.054
Avg. - - 4.192 4.596 4.561 4.614 2.908 3.176 3.208 3.101
Std. dev - - 7.369 8.053 7.915 8.109 5.483 5.287 4.796 5.144
TopGear 242 373 0.077 0.083 0.087 0.086 0.050 0.044 0.062 0.044
Bardet 120 200 0.016 0.018 0.017 0.020 0.012 0.016 0.013 0.016
Vessel 180 486 0.083 0.216 0.113 0.158 0.136 0.097 0.16 0.097
Riboflavin 71 4088 0.637 1.302 0.701 1.046 0.364 0.450 0.436 0.479
Avg. - - 0.203 0.405 0.168 0.328 0.140 0.151 0.168 0.159
Std. dev - - 0.291 0.604 0.317 0.483 0.189 0.202 0.158 0.202
Table 2: Average performance of methods across a suite of real-world datasets where the ground truth is unknown (and may not be sparse), sorted by how overdetermined the dataset is ($n/p$), and separated into the underdetermined and overdetermined cases. For both methods, the first two columns correspond to the performance of $k$-fold cross-validation, and the second two correspond to nested stability-regularized cross-validation.
We observe that for XGBoost, there is no benefit to the nested cross-validation procedure (average improvement of $-0.3\%$), and we select the same hyperparameters whether or not we account for stability via nested cross-validation on $91.5\%$ of instances. This is likely because XGBoost generates stable models by default (i.e., omitting one fold of the data hardly changes its predictions) and thus explicitly accounting for model stability does not improve its performance.
However, there is a significant benefit to nested cross-validation for CART: it improves the out-of-sample MSE by $4.1\%$ on average, with a $1.2\%$ average improvement on overdetermined datasets, and a $11.1\%$ average improvement on underdetermined datasets. Moreover, the average test set MSE is within $7.4\%$ of the average nested cross-validation error for CART (and $5.0\%$ for XGBoost), vs. $27.2\%$ inaccurate for $k$-fold cross-validation (and $8.19\%$ for XGBoost), and thus nested cross-validation also significantly reduces out-of-sample disappointment.

### v2

@@ 2.2 Generalization Bound
By combining our definitions and notation, letting $M$ represent an upper bound on the loss $\ell(y_{i},\bm{\beta}(\bm{\theta},\bm{x}_{i}))$ for any model $\bm{\beta}(\bm{\theta})$ and any data point $(\bm{x}_{i},y_{i})$ (e.g., if $(\bm{x}_{i},y_{i})$ are drawn from a bounded domain), the following result follows from Chebyshev’s inequality and upper bounds the test error $\mathbb{E}_{(x,y)\sim D}[\ell\left(y,\beta(\bm{\theta},\bm{x})\right)]$ in terms of the $k$-fold cross-validation error and the hypothesis stability (proof deferred to Section B.1):
Theorem 2.2.
@@
Suppose the training data $(\bm{x}_{i},y_{i})_{i\in[n]}$ are drawn from an unknown distribution $\mathcal{D}$ such that $M$ and $\mu_{h}$ are finite constants. Further, suppose $n$ is exactly divisible by $k$ and each $\mathcal{N}_{j}$ is of cardinality $n/k$. Then, the following bound on the test error holds with probability at least $1-\delta$ (over the draw of the training sample, and any fold-partition randomness):
$\mathbb{E}_{(x,y)\sim D}[\ell\left(y,\beta(\bm{\theta},\bm{x})\right)]$ (6)
$\leq\frac{1}{n}\sum_{j\in[k]}h_{j}(\bm{\theta})+\sqrt{\frac{M^{2}+6Mk\mu_{h}}{2k\delta}}.$
Remark 2.3 (To Train or to Validate in (6)).
@@
A similar bound to (6) can be derived using the empirical risk instead of kCV (Bousquet & Elisseeff, 2002, Theorem 11). However, this bound has a larger constant ($12$ instead of $6$) and still involves the expensive pointwise hypothesis stability defined by Bousquet & Elisseeff (2002).
@@
Theorem 2.2 reveals that, if the number of folds $k$ increases with $n$, $M$ is finite, and the hypothesis stability $\mu_{h}$ decreases with $n$, then the kCV error generalizes to the test set with high probability as $n$ becomes large. Moreover, when models have the same cross-validation error, hypothesis stability, and loss bound $M$, training on more folds results in a stronger generalization bound.
@@
We remark that Theorem 2.2 implicitly justifies the use of regularization in machine learning, because regularization implicitly controls the hypothesis stability $\mu_{h}$, leading to better generalization properties when $\mu_{h}$ is lower. Indeed, Bousquet & Elisseeff (2002, Theorem 22) provides a result formalizing this notion in the context of supervised learning for Reproducing Kernel Hilbert Spaces.
@@
Unfortunately, in preliminary experiments, we found that Equation (6)’s bound is often excessively conservative in practice, especially when $n\gg p$. This conservatism stems from using Chebyshev’s inequality in the proof of Theorem 2.2, which is known to be tight for discrete measures but excessively conservative over continuous measures (Bertsimas & Popescu, 2005), especially unimodal continuous measures (Van Parys et al., 2016). Thus, motivated by robust optimization, where probabilistic guarantees are used to motivate uncertainty sets but less stringent guarantees are used in practice to avoid excessive conservatism (see Gorissen et al., 2015, Section 3, for a discussion), we leverage Theorem 2.2 to propose a new approach to cross-validation in the rest of this section.
2.3 Stability Regularization and Nesting
@@
Motivated by the idea that more stable models are less likely to disappoint out-of-sample (Theorem 2.2), we propose selecting models that minimize a weighted sum of the kCV error and the empirical hypothesis stability. This corresponds to selecting $\bm{\theta}$ through the optimization problem
$\bm{\theta}\in\operatorname*{arg\,min}_{\bm{\theta}\in\bm{\Theta}}\frac{1}{n}\sum_{j\in[k]}h_{j}(\bm{\theta})+\lambda\hat{\mu_{h}}(\bm{\theta}).$ (7)
@@
Notably, this procedure still requires a hyperparameter $\lambda$, which the user must select. Accordingly, we invoke nested cross-validation, which has been empirically shown to be significantly less vulnerable to out-of-sample disappointment than regular cross-validation (Cawley & Talbot, 2010; Bates et al., 2021). We perform an outer loop over candidate $\lambda$ values. For each $\lambda$, we run an inner (nested) k-fold cross-validation: on each fold, train a model with that $\lambda$ (using the remaining folds with stability regularization) and measure its validation error. We then choose the $\lambda$ that yields the lowest average outer validation error across the $k$ folds.
@@
$\lambda\in$ $\arg\min_{\lambda\in\bm{\Lambda}}\sum_{j\in[k]}\sum_{i\in\mathcal{N}_{j}}\ell(y_{i},\bm{\beta}^{(\mathcal{N}_{j})}(\bm{\theta}_{j}^{\star},\bm{x}_{i})),$ (8)
where $\bm{\beta}^{(\mathcal{N}_{j})}(\bm{\theta}_{j}^{\star})$ denotes a model trained on all data but the fold $\mathcal{N}_{j}$ with hyperparameters $\bm{\theta}_{j}^{\star}$, and $\bm{\theta}_{j}^{\star}$ denotes an optimal solution to the following lower-level problem, which cross-validates $\bm{\theta}$ with the $j$th fold of the data omitted and $\lambda$ fixed, so that $\bm{\theta}^{\star}_{j}$ is in the argmin of
$\min_{\bm{\theta}\in\bm{\Theta}}\frac{1}{n-|\mathcal{N}_{j}|}\sum_{l\in[k]:l\neq j}\sum_{i\in\mathcal{N}_{l}}\ell(y_{i},\bm{\beta}^{(\mathcal{N}_{j,l})}(\bm{\theta}))+\lambda\hat{\mu}_{h}^{(-j)}(\bm{\theta}),$
where $\hat{\mu}_{h}^{(-j)}(\bm{\theta})$ is calculated without reference to the $j$th fold of the data, $\mathcal{N}_{j}$, in this case. This corresponds to selecting $\lambda$ in a manner that ensures that models $\bm{\theta^{\star}_{j}}$ perform well on average, as in the standard nested cross-validation paradigm.
Finally, once $\lambda=\lambda^{\star}$ is fixed, we select $\bm{\theta}$ by minimizing (7). This selects models that are robust to omitting one fold of the data and tend to perform well out-of-sample, as estimated by nested cross-validation, at the cost of increased hyperparameter selection time. We formalize this procedure in Algorithm 1 (see Appendix C).
Complexity analysis: In terms of the complexity of Algorithm 1, let $|\Theta|$ be the number of hyperparameter candidates, and let $|\Lambda|$ be the number of candidate stability weights. Then, standard $k$-fold cross-validation (kCV) evaluates each hyperparameter combination $\theta$ using $k$ model fits, for a total of $O(k|\Theta|)$ model calls, plus one at the selected model. On the other hand, Algorithm 1 adds an outer $k$-fold loop to tune $\lambda$; for each $\lambda$, each outer fold $t$, and each evaluated $\theta$, the inner loop fits 1+(k1)=k1+(k-1)=k models to compute both the inner CV score and the stability proxy. Thus, a direct implementation of Algorithm 1 requires $O(|\Lambda|k^{2}|\Theta|)$ model fits to select $\lambda^{\star}$, plus $O((k+1)|\Theta|)$ additional fits to select $\theta^{\star}$ at the chosen $\lambda^{\star}$. Relative to standard kCV, this is an overhead of order $|\Lambda|k$ in the number of fitted models. In the special case where the inner search exhaustively evaluates a fixed finite grid $\Theta$, this can be reduced by a factor of $|\Lambda|$ (for a total overhead factor of $k$) by saving all models and computing the optimal $\lambda$ after one pass of the inner loop. However, for adaptive searches (where the set of evaluated $\theta$ depends on $\lambda$, as occurs in our numerics), the $O(|\Lambda|k^{2}|\Theta|)$ bound captures the worst-case fit count.
@@ 3 Numerical Experiments
In this section, we evaluate the numerical performance of the nested regularized cross-validation scheme proposed in Section 2. All experiments were implemented in Julia version $1.9$, using Mosek version $11.0$ to solve all conic optimization problems, and conducted on a MacBook Pro with Apple M3 cores and $36$ GB RAM. Where needed, we used the RCall package to access R version $4.2.1$, which calls certain baseline methods implemented in R, via Julia.
@@ 3.1 Ridge Regularized Best Subset Selection
We now benchmark our proposed nested cross-validation scheme on sparse regression for a suite of commonly studied real-world datasets. Specifically, we benchmark a cyclic coordinate descent scheme for $\ell_{0}$$\ell_{2}^{2}$ sparse regression, where we repeatedly solve the following lower-level problem [^1] Note that the terms $\beta_{j}^{2}/z_{j}$ implicitly impose the logical constraints $\beta_{j}=0$ if $z_{j}=0$ using the perspective reformulation trick (Günlük & Linderoth, 2010). for different values of the sparsity parameter $\tau$ and the regularization hyperparameter $\gamma$
$\min_{\bm{\beta}\in\mathbb{R}^{p},\bm{z}\in\{0,1\}^{p}}\;$ $\frac{1}{2\gamma}\sum_{j=1}^{p}\frac{\beta_{j}^{2}}{z_{j}}+\frac{1}{2}\|\bm{X}\bm{\beta}-\bm{y}\|_{2}^{2}\ \text{s.t.}\ \sum_{j=1}^{p}z_{j}\leq\tau,$
using the greedy rounding algorithm described by (Bertsimas et al., 2021) to obtain near-optimal solutions in a practically tractable amount of time. In particular, we iteratively minimize the $k$-fold cross-validation error with respect to $\tau$ (with $\gamma$ fixed) and with respect to $\gamma$ (with $\tau$ fixed) until we either cycle or exceed a limit of $10$ iterations. After converging, we fit a model to the full dataset with the hyperparameters ($\tau^{\star},n_{\text{train}}\gamma^{\star}/n)$, where $n_{\text{train}}$ is the number of observations with one fold of the data left out as in (Liu & Dobriban, 2020), to account for the extra fold when fitting a model to the full dataset. Moreover, we perform the same procedure for nested $k$-fold cross-validation with stability regularization. Note that for our cyclic coordinate descent schemes, we set the largest permissible value of $\tau$ such that $\tau\log\tau\leq n$, because Gamarnik & Zadik (2022, Theorem 2.5) demonstrated that, up to constant terms and under certain assumptions on the data generation process, on the order of $\tau\log\tau$ observations are necessary to recover a sparse model with binary coefficients. In preliminary experiments, we relaxed this requirement to $\tau\leq p$ and found that this did not change the optimal value of $\tau$. Moreover, we use a grid size of $20$ values of $\gamma$ log-uniformly distributed over $10^{-3},10^{3}]$, and $10$ values of $\lambda$ log-uniformly distributed over $10^{-4},10^{4}]$.
@@
  • The L0Learn.cvfit method implemented in the L0Learn R package (cf. Hazimeh & Mazumder, 2020), with a grid of $10$ different values of $\gamma$ and default parameters otherwise.
@@
Table 1 depicts the dimensionality of each dataset, the average $k$-fold cross-validation error (“CV”) or nested $k$-fold cross-validation error (“nCV”), the average test set error (“MSE”), and the sparsity attained by our coordinate descent scheme without any stability regularization or nesting (kCV), our cyclic coordinate descent with nested stability-regularized cross-validation (nested-kCV), and the performance of MCP on each dataset. We remark that the average (geometric mean over the average for each dataset) runtime was $14.8$ seconds for sparse regression, $592$ seconds for sparse regression with nesting, and under $1$ second for all other methods. Table 4 (Appendix E) also shows the performance of other methods from the literature (SCAD, GLMNet, and L0Learn) in the same datasets.
We measure the improvement from nested cross-validation vs. $k$-fold cross-validation by computing the average MSE ratio for each dataset $d$, $r_{d}:=\text{MSE}_{nested,d}/\text{MSE}_{kCV,d}$, then aggregating by taking the geometric mean of $r_{d}$ to compute an average improvement (accounting for the fact that percentage improvement is an asymmetric measure).
@@
Table 1: Average performance of methods across a suite of real-world datasets where the ground truth is unknown (and may not be sparse), sorted by how overdetermined the dataset is ($n/p$), and separated into the underdetermined and overdetermined cases.
We observe that across the overdetermined datasets, nested cross-validation improves the out-of-sample MSE of sparse regression by $10.0\%$ on average, while across the underdetermined datasets it worsens the out-of-sample MSE by $5.92\%$ on average, for an overall improvement of $4.85\%$ on average. Notably, the hyperparameters selected by both methods agree $67\%$ of the time, meaning the average improvement conditioned on instances selected when the hyperparameters selected by both methods disagree is above $10\%$. Moreover, the nested kCV error is, on average, $0.9\%$ smaller than the test set error for the stability-regularized method, while the kCV error is on average $21.8\%$ smaller than the test set error for sparse ridge regression. This is especially visible for the most underdetermined datasets (Vessel and Riboflavin), where the kCV error for sparse ridge regression is lower than all other methods, but this does not translate to better out-of-sample performance. In contrast, the nested kCV error is a substantially more accurate estimator of the test set error.
@@ 3.2 Tree-Based Methods
We benchmark cyclic coordinate descent for the following two methods, using $10$ values of $\lambda$ log-uniformly distributed over $10^{-4},10^{4}]$ for our nesting procedure:
  • A Julia implementation of CART (Breiman et al., 1984) via the DecisionTree.jl package, where we iteratively optimize the $5$-fold and nested $5$-fold cross-validation error with respect to the tree depth over a grid of the integers $\{1,\ldots,10\}$ and the min_samples parameter over a grid of the integers $\{2,\ldots,10\}$, with the tree depth initially fixed to $5$ for both approaches. Note that the DecisionTree.jl package holds all unspecified parameters to their default values, and thus all remaining CART parameters will take default parameters in this experiment.
  • A Julia implementation of XGBoost (Chen & Guestrin, 2016) via the XGBoost.jl package, where we iteratively optimize the $5$-fold and nested $5$-fold cross-validation error with respect to the max_depth parameter over a grid of the integers $\{1,\ldots,10\}$ and the subsample parameter over the grid $\{0.01,0.02,0.03,\ldots,1.0\}$, with the tree depth initially fixed to $5$ for both approaches.
Table 2 depicts the dimensionality of each dataset, the average $k$-fold cross-validation error (“CV”) or nested $k$-fold cross-validation error (“nCV”), and the average test set error (“MSE”), with and without nested cross-validation. For both methods, the first two columns correspond to $k$-fold cross-validation, and the last two columns to nested stability-regularized cross-validation. Note that we repeat our train/test split procedure 50 times for each dataset and report the average to reduce the variance in our results.
As in the previous experiment, we measure the improvement from nested cross-validation vs. $k$-fold cross-validation by computing the average MSE across each dataset, normalizing by the MSE of $k$-fold cross-validation on that dataset to compute a percentage improvement across each dataset, and then taking the geometric mean across all datasets (to account for the fact that percentage improvement is an asymmetric measure). The average (geometric mean over averages for each dataset) runtime was $8.07$s (resp. $377.1$s) for XGBoost without (with) nesting, and <0.01<0.01s (resp. <0.01<0.01s) for CART with/without nesting.
@@
Wine 6497 11 0.534 0.531 0.539 0.531 0.422 0.405 0.429 0.405
Housing 506 13 17.385 19.183 18.912 19.040 11.092 12.464 12.155 12.464
AutoMPG 392 25 15.351 15.878 16.55 16.004 10.36 11.872 11.395 11.628
Hitters 263 19 0.050 0.057 0.055 0.058 0.038 0.039 0.042 0.039
Prostate 97 8 0.717 0.788 0.815 0.796 0.581 0.696 0.648 0.679
Servo 167 19 0.520 0.565 0.656 0.565 0.237 0.237 0.351 0.237
Toxicity 38 9 0.081 0.118 0.092 0.114 0.066 0.081 0.078 0.081
Steam 25 8 1.089 1.308 1.47 1.244 0.91 1.064 1.165 1.118
Alcohol2 44 21 0.850 1.110 0.915 1.071 1.123 1.264 1.289 1.253
TopGear 242 373 0.076 0.087 0.086 0.086 0.051 0.052 0.061 0.052
Bardet 120 200 0.015 0.020 0.017 0.020 0.012 0.014 0.013 0.014
Vessel 180 486 0.063 0.149 0.084 0.136 0.111 0.095 0.136 0.095
Riboflavin 71 4088 0.719 0.979 0.817 0.904 0.404 0.463 0.496 0.465
Table 2: Average performance of methods across a suite of real-world datasets where the ground truth is unknown (and may not be sparse), sorted by how overdetermined the dataset is ($n/p$), and separated into the underdetermined and overdetermined cases. For both methods, the first two columns correspond to $k$-fold cross-validation, and the last two columns to nested stability-regularized cross-validation.
We observe that for XGBoost, there is no benefit to the nested cross-validation procedure (average improvement of $0.6\%$), and we select the same hyperparameters whether or not we account for stability via nested cross-validation on $92\%$ of instances. This is likely because XGBoost generates stable models by default (i.e., omitting one fold of the data hardly changes its predictions), so explicitly accounting for model stability does not improve performance.
However, there is a benefit to nested cross-validation for CART: it improves the out-of-sample MSE by $2.2\%$ on average, with a $1.1\%$ average improvement on overdetermined datasets, and a $4.8\%$ average improvement on underdetermined datasets. Because both methods selected the same hyperparameters on $51\%$ of all instances, this corresponds to an average improvement of over $4\%$ on instances where the two methods did not select the same hyperparameters. Moreover, the average test set MSE is within $6.2\%$ of the average nested cross-validation error for CART (and $7.4\%$ for XGBoost), vs. $24.7\%$ inaccurate for $k$-fold cross-validation (and $8.1\%$ for XGBoost), and thus nested cross-validation also reduces out-of-sample disappointment.